# Exact path integrals on half-line in quantum cosmology with a fluid clock and aspects of operator ordering ambiguity

### v1

@@ Abstract
We perform exact half-line path integral quantization of flat, homogeneous cosmological models containing a perfect fluid acting as an internal clock, in a $D+1$ dimensional minisuperspace setup. We also discuss certain classes of operator ordering ambiguity inherent in such quantization procedures and argue that a particular ordering prescription in the quantum theory can preserve two symmetries, namely arbitrary lapse rescalings and general covariance, which are already present at the classical level. As a result of this imposition, a large class of quantum Hamiltonians differing by operator ordering produces the same inner products between quantum states. This imposition of the two symmetries of the classical minisuperspace models leads to a unique prescription for writing the quantum Hamiltonian for minisuperspace dimension $D\geq 2$. Interestingly, in the case of $D=1$, the lapse rescaling symmetry is lost in the quantum theory, leading to an essentially ambiguous description of the canonical theory. We provide general proof of this in the context of both canonical quantization and path integrals. We supply concrete examples to validate our findings further.
@@ I Introduction
Our central findings are as follows: it can be shown within both canonical and path integral formalisms that large classes of operator ordering ambiguities of the quantum Hamiltonian constraint do not affect the physical predictions of the theory—such as the inner product between states, quantum correlators, probability measure, etc.—if (a) there is a reference fluid working as the intrinsic clock degree of freedom describing the unitary evolution of the quantum state of the universe, (b) apart from the clock and the scale factor of the universe, there is at least one additional dynamical degree of freedom in the minisuperspace, such as a scalar field, that is, the dimension of the minisuperspace (excluding the clock) is greater than or equal to $2$, (c) the quantum Hamiltonian is written in a generally covariant form (the Laplace-Beltrami form), and (d) the potential receives a unique and well-defined quantum correction of the order $\mathcal{O}(\hbar^{2})$ depending on the underlying geometry of the minisuperspace through the Ricci curvature scalar such that the quantum Hamiltonian constraint is preserved under arbitrary lapse rescalings. We provide proof of these results in both canonical and path integral formalisms.
@@
    Due to Hamiltonian constraint for a generally covariant system, quantum gravity has a problem of time, that is, the “Scrödinger equation” for quantum gravity, which is called the Wheeler-DeWitt equation is timeless—in explicit notation, this reads as ^$\hat{\mathscr{H}}\Psi=0$ with no extrinsic time evolution. In the absence of time, it then becomes difficult to describe the unitary evolution of the universe with respect to time or even properly define the Hilbert space of states and interpret probabilities. Among many proposed solutions [44], we consider the relational approach to the problem of time, wherein a perfect fluid is introduced as a reference fluid, which acts as the clock degree of freedom [45, 46, 47, 48]. The fluid acts as an internal clock for the system against which the unitary evolution of quantum states of the universe can be described.
@@
In what follows, we write down the general Wheeler-DeWitt equation for a $D+1$ dimensional minisuperspace, where $D$ denotes the dimension of the minisuperspace without any clock, and $+1$ corresponds to the reference fluid acting as a clock. We then classify the possible ambiguities in writing the quantum Hamiltonian operator into three categories. We show, following [7, 8], that if a particular class of the ambiguity that appears as an order $\mathcal{O}(\hbar^{2})$ correction to the superpotential in the Wheeler-DeWitt equation is chosen to be proportional to the Ricci curvature scalar of the minisuperspace (with the proportionality constant being $+\frac{(D-2)}{8(D-1)}$), then the quantum theory becomes conformally invariant in the sense that the inner product and the quantum correlators do not get affected by arbitrary lapse function rescalings. When conformal and general covariance, which are symmetries of the classical theory, are both present in the quantum theory as well then (for $D\geq 2$) the other classes of the ambiguities in the quantum Hamiltonian do not affect the physical predictions. Following the proof in the canonical formalism, we also present proof of this result in the path integral approach. For the case of $D=1$, these symmetry-guided arguments fail and operator ambiguities become an essential feature of the quantum theory. Our work is different from [7, 8] in two key ways: that we have an additional clock in the system which allows us to define unitary evolution of the quantum states and positive definite probability measure, and we also extend the proofs in the context of path integral as well.
@@ II.1 Canonical quantum gravity and Wheeler-DeWitt equation
where we have used the identification $p_{T}\to-i\hbar\frac{\partial}{\partial T}$, and $\mathscr{H}\to\hat{\mathscr{H}}$. Due to the linear dependence of the Hamiltonian of the fluid on its momentum, the fluid variable $T$ now acts as the time parameter for the evolution of the wave function of the universe. The precise form of the quantum Hamiltonian operator $\hat{\mathscr{H}}$ depends on the system and it generates the translations in the fluid variable $T$. Therefore, ensuring the self-adjointness of this Hamiltonian implies a unitary evolution in the fluid clock. Moreover, as mentioned before, the form of the operator $\hat{\mathscr{H}}$ can have ordering ambiguities and we discuss this issue in the following.
@@ II.2 Breaking of the lapse rescaling symmetry as operator ordering ambiguity
From the action 2 it is clear that the classical minisuperspace model has two very crucial symmetries—general covariance with respect to the minisuperspace metric [^1] The general covariance is under the arbitrary (minisuperspace) coordinate transformations $q^{A}\to\tilde{q}^{A}(q^{B})$. These transformations are restricted not to violate the assumption of homogeneity of the universe. and conformal invariance under arbitrary lapse rescalings $N\to\tilde{N}\Omega^{-2}$. Notice that under the lapse rescaling, the classical Hamiltonian changes by an overall factor, and the classical Hamiltonian constraint remains unchanged. This lapse rescaling is equivalent to time reparameterization ${\rm d}t\to\Omega^{-2}{\rm d}t$, and thus the symmetry under lapse rescaling is equivalent to the time reparametrization invariance.
@@
where we have defined $\tilde{\mathscr{G}}^{AB}=\Omega^{-2}\mathscr{G}^{AB}$, and $\tilde{U}=\Omega^{-2}U$. As the classical theory remains unchanged by such rescalings, this new action is as good as the previous one in 2. We could in principle start to quantize this action in 8 instead of the one in 2. However, the resulting Wheeler-DeWitt equation for the new theory is
@@ II.3 Conformal invariance and resolution of the operator ordering ambiguity
Notice that the invariance of the inner products and the probability measure depended on the fact that under the lapse rescaling, the integration measure changes not by $\sqrt{|\mathscr{G}|}\to\sqrt{|\tilde{\mathscr{G}}|}$, but by an additional factor of $\Omega^{-2}$, that is, correctly the measure transforms as $\sqrt{|\mathscr{G}|}\to\sqrt{|\tilde{\mathscr{G}}|}\Omega^{-2}$. This additional factor appears due to the transformation $\partial/\partial T\to\Omega^{-2}\partial/\partial T$, and is lost in the case of the timeless formalisms to quantum cosmology. Thus, having a clock in the theory is crucial to ensure the fact that the inner products or the physical predictions of the theory are ambiguity-free and do not depend on the operator ordering choices (of conformal-class) of the quantum Hamiltonian.
@@ III Path integral quantization: general formalism
The minisuperspace model we are dealing with is a constrained system, and as a result, defining the path integral kernel has to be done carefully. The constraints correspond to gauge symmetries in the system, which leads to overcounting of the paths (or histories) that are equivalent due to the symmetry. Therefore, constructing the path integral is quite non-trivial. The task of constructing the path integral in the context of gravitational systems was taken upon by Tietelboim (presently Bunster) [29, 30, 31, 32], Barvinsky and Ponomariov [35, 36], Kuchař and Hartle [73, 33, 34]. However, all of these works dealt with the problem only at a formal level. Later, Halliwell [7] explicitly constructed the path integral for the minisuperspace cosmological models (also see Halliwell and Hartle [37]), and that framework will be the basis for the discussions in the current paper. Halliwell showed that the imposition of the gauge fixing condition $\dot{N}=\chi(a,p_{a},N)$ can be done using the BFV (Batalin-Fradkin-Vilkovisky) formalism in which one adds to the gauge fixed action additional anti-commuting ghost degrees of freedom such that the BRST (Becchi-Rouet-Stora-Tyutin) symmetry is respected. Due to the BRST symmetry, the path integral will not depend on the choice of the function $\chi$. Even though [7] does not consider the additional fluid $(T,p_{T})$, the arguments presented in that work can be straightforwardly generalized to the present case. In the proper time gauge $\dot{N}=0$, the Feynman propagator has the form
@@ III.2 Conformal invariance under lapse rescaling
In the last line, we have used the integral representation of the Dirac delta function, and the convention for the normalization of the delta function has been done with a symmetric combination of both initial and final scale factors, which is consistent with the product form definition of the path integral. Recalling the discrete definition of the path integral III, we see that the time transformation 41 has to be defined in a symmetric manner over the pre and post points, that is, $(\Omega^{(j)})^{-1}(\Omega^{(j-1)})^{-1}\Delta t=\Delta u^{(j)}$. Where $\Delta t=\epsilon$ is the discrete time interval that remains uniform throughout the lattice. On the other hand, the new time interval $\Delta u^{(j)}$ is not uniform and itself becomes time-dependent. The overall factor $\Omega^{-2}$ with the Hamiltonian in III.2 can also be written as $(\Omega^{(j)}\Omega^{(j-1)})^{-1}$, since the difference of this quantity from $(\Omega^{(j)})^{-2}$ appears in the path integral at a greater order than $\mathcal{O}(\epsilon)$ and can be neglected in the limit $\epsilon\to 0$. In the following, we continue to use the continuous version of the expressions.
@@
is nonzero only if $\mathscr{T}>0$, hence the result is a Heaviside theta function $\Theta(\mathscr{T})$. Then, the Feynman propagator is only nonzero if $\mathscr{T}=T_{\sf f}-T_{\sf i}>0$. Therefore, the clock degree of freedom has introduced the notion of causal propagation and introduced an arrow of time.
@@
To summarize, building quantum theories demanding general covariance under point canonical transformations and conformal invariance under the lapse rescaling (in $D\geq 2$) leads to unique physical predictions for a large class of operator-ordering ambiguous quantum Hamiltonians. In $D=1$ minisuperspace, the symmetry-guided arguments constructed above do not work, and there is an ambiguity associated with lapse rescaling, which we will show in the next section, with an explicit example, is related to the operator ordering ambiguity.
@@ IV.2.2 Quantization
is independent of the ambiguity parameter $p$. Unlike the case of $D=1$, the index of the Bessel function, here, does not depend on the parameter $p$. The parameter $p$ appears in the wave function only as an overall factor $a^{-(1+p)}$, and gets canceled by the integration measure in the inner product. In Appendix C, we show that the expectation value of the geometric observables is also independent of the ambiguity parameter, implying that the quantum cosmology in $D\geq 2$ or in this particular example $D=4$, is ambiguity-free: infinitely many different quantum Hamiltonians parametrized by $p$ (conformal-class ambiguous Hamiltonians) have been rendered equivalent by a simple choice of the potential-class ambiguity term.
@@ V Examples of exactly solvable cases: path integral formalism
In the following, we shall discuss the exact path integral quantization of two flat, homogeneous cosmologies: (a) the isotropic FLRW, and (b) the anisotropic Bianchi I spacetime, along with the action for a homogeneous, massless, free scalar field. The case (a) will constitute an example of the quantization of minisuperspace of dimension $D=1$. In this case, after the quantization, the lapse rescaling symmetry is lost and we shall show that we can recover the solutions corresponding to a one-parameter family of Wheeler-Dewitt equations. For conformally invariant quantum theory we need to have a minisuperspace of dimension $D\geq 2$. The case (b) will constitute an example of $D=4$ minisuperspace. We shall show that, in this case, the wave function conformally transforms under lapse rescalings.

### v2

@@ Abstract
We perform exact half-line path integral quantization of flat, homogeneous cosmological models containing a perfect fluid acting as an internal clock, in a $D+1$ dimensional minisuperspace setup. We also discuss certain classes of operator ordering ambiguity inherent in such quantization procedures and argue that a particular ordering prescription in the quantum theory can preserve two symmetries, namely arbitrary lapse rescalings and general covariance, which are already present at the classical level. As a result of this imposition, a large class of quantum Hamiltonians differing by operator ordering produces the same inner products between quantum states. This imposition of the two symmetries of the classical minisuperspace models leads to a unique prescription for writing the quantum Hamiltonian for minisuperspace dimension $D>2$. Interestingly, in the case of $D=1$, the lapse rescaling symmetry is lost in the quantum theory, leading to an essentially ambiguous description of the canonical theory. We provide general proof of this in the context of both canonical quantization and path integrals. We supply concrete examples to validate our findings further.
@@ I Introduction
Our central findings are as follows: it can be shown within both canonical and path integral formalisms that large classes of operator ordering ambiguities of the quantum Hamiltonian constraint do not affect the physical predictions of the theory—such as the inner product between states, quantum correlators, probability measure, etc.—if (a) there is a reference fluid working as the intrinsic clock degree of freedom describing the unitary evolution of the quantum state of the universe, (b) apart from the clock and the scale factor of the universe, there is at least two additional dynamical degree of freedom in the minisuperspace, such as a scalar field, that is, the dimension of the minisuperspace (excluding the clock) is greater than $2$, (c) the quantum Hamiltonian is written in a generally covariant form (the Laplace-Beltrami form), and (d) the potential receives a unique and well-defined quantum correction of the order $\mathcal{O}(\hbar^{2})$ depending on the underlying geometry of the minisuperspace through the Ricci curvature scalar such that the quantum Hamiltonian constraint is preserved under arbitrary lapse rescalings. We provide proof of these results in both canonical and path integral formalisms.
@@
    Due to Hamiltonian constraint for a generally covariant system, quantum gravity has a problem of time, that is, the “Scrödinger equation” for quantum gravity, which is called the Wheeler-DeWitt equation is timeless—in explicit notation, this reads as ^$\hat{\mathscr{H}}\Psi=0$ with no extrinsic time evolution. In the absence of time, it then becomes difficult to describe the unitary evolution of the universe with respect to time or even properly define the Hilbert space of states and interpret probabilities. Among many proposed solutions [44], we consider the relational approach to the problem of time, wherein a perfect fluid is introduced as a reference fluid, which acts as the clock degree of freedom [45, 46, 47, 48]. The fluid acts as an internal clock for the system against which the unitary evolution of quantum states of the universe can be described and this framework can be used to study early universe physics [49, 50, 51, 52, 53, 54, 55, 56, 57].
@@
In what follows, we write down the general Wheeler-DeWitt equation for a $D+1$ dimensional minisuperspace, where $D$ denotes the dimension of the minisuperspace without any clock, and $+1$ corresponds to the reference fluid acting as a clock. We then classify the possible ambiguities in writing the quantum Hamiltonian operator into three categories. We show, following [7, 8], that if a particular class of the ambiguity that appears as an order $\mathcal{O}(\hbar^{2})$ correction to the superpotential in the Wheeler-DeWitt equation is chosen to be proportional to the Ricci curvature scalar of the minisuperspace (with the proportionality constant being $+\frac{(D-2)}{8(D-1)}$), then the quantum theory becomes conformally invariant in the sense that the inner product and the quantum correlators do not get affected by arbitrary lapse function rescalings. When conformal and general covariance, which are symmetries of the classical theory, are both present in the quantum theory as well then (for $D>2$) the other classes of the ambiguities in the quantum Hamiltonian do not affect the physical predictions. Following the proof in the canonical formalism, we also present proof of this result in the path integral approach. For the case of $D=1$, these symmetry-guided arguments fail and operator ambiguities become an essential feature of the quantum theory. Our work is different from [7, 8] in two key ways: that we have an additional clock in the system which allows us to define unitary evolution of the quantum states and positive definite probability measure, and we also extend the proofs in the context of path integral as well.
@@ II.1 Canonical quantum gravity and Wheeler-DeWitt equation
where we have used the identification $p_{T}\to-i\hbar\frac{\partial}{\partial T}$, and $\mathscr{H}\to\hat{\mathscr{H}}$. Due to the linear dependence of the Hamiltonian of the fluid on its momentum, the fluid variable $T$ now acts as the time parameter for the evolution of the wave function of the universe. The precise form of the quantum Hamiltonian operator $\hat{\mathscr{H}}$ depends on the system, and it generates the translations in the fluid variable $T$. Therefore, ensuring the self-adjointness of this Hamiltonian implies a unitary evolution with respect to the fluid clock. Moreover, as mentioned before, the form of the operator $\hat{\mathscr{H}}$ can have ordering ambiguities, and we discuss this issue in the following.
We note that ‘fluid’ is an effective description of matter, and at a more fundamental level, all matter should be described as renormalizable quantum fields. Therefore, the above treatment works insofar as the Schutz fluid description (see appendix A) is valid, which may break at high energies. However, we would like to emphasize that the precise nature of the fluid description is not essential for our treatment. What is required is that we have a degree of freedom whose Hamiltonian can be brought to have a linear dependence on its momentum using canonical transformation, such that upon quantization, this degree of freedom can play the role of a clock. A perfect fluid is an ideal demonstration of this principle; however, by no means is it a unique choice for the clock, and other degrees of freedom may play a similar role. The formalism discussed here applies to any degree of freedom with Hamiltonian depending linearly on the momentum. Furthermore, in minisuperspace approximation, it is always assumed that relevant energy scales are much below the Planck scale so that arbitrary metric fluctuations (e.g. inhomogeneities) can be ignored or treated perturbatively.
@@ II.2 Breaking of the lapse rescaling symmetry as operator ordering ambiguity
From the action 2 it is clear that the classical minisuperspace model has two very crucial symmetries—general covariance with respect to the minisuperspace metric [^1] The general covariance is under the arbitrary (minisuperspace) coordinate transformations $q^{A}\to\tilde{q}^{A}(q^{B})$. These transformations are restricted so as not to violate the assumption of homogeneity of the universe. and conformal invariance under arbitrary lapse rescalings $N\to\tilde{N}\Omega^{-2}$. Notice that under the lapse rescaling, the classical Hamiltonian changes by an overall factor, and the classical Hamiltonian constraint remains unchanged. This lapse rescaling is equivalent to time reparameterization ${\rm d}t\to\Omega^{-2}{\rm d}t$, and thus the symmetry under lapse rescaling is equivalent to the time reparametrization invariance. Note that the factor $\Omega(\bm{q},\bm{p})$ can, in principle, be any arbitrary non-vanishing function of the phase space since recall from the form of the action in 2, that we are working with the first order form of the action which depends on both $\bm{q}$ and $\bm{p}$. However, in the following, we restrict ourselves to the case where $\Omega(\bm{q})$ is only a function of the configuration space since this form will be useful for our discussions on the operator ordering ambiguities.
@@
where we have defined $\tilde{\mathscr{G}}^{AB}=\Omega^{-2}\mathscr{G}^{AB}$, and $\tilde{U}=\Omega^{-2}U$. Here, notice that restricting $\Omega(\bm{q})$ to be only a function of the configuration space (and not the full phase space) allows us to reinterpret this lapse rescaling as the Weyl transformation of the minisuperspace metric. These transformations are less general than the time reparametrization invariance, which is characterized by $\Omega(\bm{q},\bm{p})$, a function of the full phase space. As the classical theory remains unchanged by such rescalings, this new action is as good as the previous one in 2. We could in principle start to quantize this action in 8 instead of the one in 2. However, the resulting Wheeler-DeWitt equation for the new theory is
@@
The fact that $\Omega^{D-2}$ can be identified with $F_{1}$ works when $D\neq 2$. Therefore, in $D=2$, building a quantum theory that is conformally invariant (this requires the choice $\xi=0$) does not give any handle on the ambiguity functions $F_{1}$ and, therefore, the operator ordering ambiguity cannot be tackled based on symmetries.
@@ II.3 Conformal invariance and resolution of the operator ordering ambiguity
Notice that the invariance of the inner products and the probability measure depended on the fact that under the lapse rescaling, the integration measure changes not by $\sqrt{|\mathscr{G}|}\to\sqrt{|\tilde{\mathscr{G}}|}$, but by an additional factor of $\Omega^{-2}$, that is, correctly the measure transforms as $\sqrt{|\mathscr{G}|}\to\sqrt{|\tilde{\mathscr{G}}|}\Omega^{-2}$. This additional factor appears due to the transformation $\partial/\partial T\to\Omega^{-2}\partial/\partial T$, and is lost in the case of the timeless formalisms to quantum cosmology. Thus, having a clock in the theory is crucial to ensure the fact that the inner products or the physical predictions of the theory are ambiguity-free and do not depend on the operator ordering choices (of conformal-class) of the quantum Hamiltonian. We note that it is possible to define a probability measure that respects the conformal invariance by construction, even in the absence of a clock. For example, in [79] the probability measure $|\Psi|^{\frac{2D}{D-2}}\,\,{\rm dvol}=|\Psi|^{\frac{2D}{D-2}}\sqrt{|g|}{\rm d}^{D}q$ is conformally invariant by construction. However, in our case, we did not construct the probability measure by demanding conformal invariance. We simply used the Sturm-Liouville procedure to determine the weight factor in the usual inner product in non-relativistic quantum mechanics, demanding only the self-adjointness of the Hamiltonian operator. Surprisingly, this natural inner product also turns out to be a conformally invariant inner product as well.
@@ III Path integral quantization: general formalism
The minisuperspace model we are dealing with is a constrained system, and as a result, defining the path integral kernel has to be done carefully. The constraints correspond to gauge symmetries in the system, which leads to overcounting of the paths (or histories) that are equivalent due to the symmetry. Therefore, constructing the path integral is quite non-trivial. The task of constructing the path integral in the context of gravitational systems was taken upon by Teitelboim (presently Bunster) [29, 30, 31, 32], Barvinsky and Ponomariov [35, 36], Kuchař and Hartle [83, 33, 34]. However, all of these works dealt with the problem only at a formal level. Later, Halliwell [7] explicitly constructed the path integral for the minisuperspace cosmological models (also see Halliwell and Hartle [37]), and that framework will be the basis for the discussions in the current paper. Halliwell showed that the imposition of the gauge fixing condition $\dot{N}=\chi(a,p_{a},N)$ can be done using the BFV (Batalin-Fradkin-Vilkovisky) formalism in which one adds to the gauge fixed action additional anti-commuting ghost degrees of freedom such that the BRST (Becchi-Rouet-Stora-Tyutin) symmetry is respected. Due to the BRST symmetry, the path integral will not depend on the choice of the function $\chi$. Even though [7] does not consider the additional fluid $(T,p_{T})$, the arguments presented in that work can be straightforwardly generalized to the present case. In the proper time gauge $\dot{N}=0$, the Feynman propagator has the form
@@ III.2 Conformal invariance under lapse rescaling
In the last line, we have used the integral representation of the Dirac delta function, and the convention for the normalization of the delta function has been done with a symmetric combination of both initial and final scale factors, which is consistent with the product form definition of the path integral.
Recalling the discrete definition of the path integral III, we see that the time transformation 41 has to be defined in a symmetric manner over the pre and post points, that is, $(\Omega^{(j)})^{-1}(\Omega^{(j-1)})^{-1}\Delta t=\Delta u^{(j)}$. Where $\Delta t=\epsilon$ is the discrete time interval that remains uniform throughout the lattice. On the other hand, the new time interval $\Delta u^{(j)}$ is not uniform and itself becomes time-dependent. The overall factor $\Omega^{-2}$ with the Hamiltonian in III.2 can also be written as $(\Omega^{(j)}\Omega^{(j-1)})^{-1}$, since the difference of this quantity from $(\Omega^{(j)})^{-2}$ appears in the path integral at a greater order than $\mathcal{O}(\epsilon)$ and can be neglected in the limit $\epsilon\to 0$. In the following, we continue to use the continuous version of the expressions.
@@
is nonzero only if $\mathscr{T}>0$, hence the result is a Heaviside theta function $\Theta(\mathscr{T})$. Then, the Feynman propagator is only nonzero if $\mathscr{T}=T_{\sf f}-T_{\sf i}>0$, i.e., it is now a causal propagator. As a result, the fluid variable $T$ can be interpreted as a meaningful clock and we have an arrow of time. Notice that this is a direct result of the restriction of the range of $u_{\sf f}$ to the half-line in Eq. III.2. Had we considered a full-line integration of $u_{\sf f}$, the result would not have been a causal propagator, justifying our choice for the range of integration.
@@
To summarize, building quantum theories demanding general covariance under point canonical transformations and conformal invariance under the lapse rescaling (in $D>2$) leads to unique physical predictions for a large class of operator-ordering ambiguous quantum Hamiltonians. In $D=1$ minisuperspace, the symmetry-guided arguments constructed above do not work, and there is an ambiguity associated with lapse rescaling, which we will show in the next section, with an explicit example, is related to the operator ordering ambiguity.
@@ IV.2.2 Quantization
is independent of the ambiguity parameter $p$. Unlike the case of $D=1$, the index of the Bessel function, here, does not depend on the parameter $p$. The parameter $p$ appears in the wave function only as an overall factor $a^{-(1+p)}$, and gets canceled by the integration measure in the inner product. In Appendix C, we show that the expectation value of the geometric observables is also independent of the ambiguity parameter, implying that the quantum cosmology in $D>2$ or in this particular example $D=4$, is ambiguity-free: infinitely many different quantum Hamiltonians parametrized by $p$ (conformal-class ambiguous Hamiltonians) have been rendered equivalent by a simple choice of the potential-class ambiguity term.
@@ V Examples of exactly solvable cases: path integral formalism
In the following, we shall discuss the exact path integral quantization of two flat, homogeneous cosmologies: (a) the isotropic FLRW, and (b) the anisotropic Bianchi I spacetime, along with the action for a homogeneous, massless, free scalar field. The case (a) will constitute an example of the quantization of minisuperspace of dimension $D=1$. In this case, after the quantization, the lapse rescaling symmetry is lost and we shall show that we can recover the solutions corresponding to a one-parameter family of Wheeler-Dewitt equations. For conformally invariant quantum theory we need to have a minisuperspace of dimension $D>2$. The case (b) will constitute an example of $D=4$ minisuperspace. We shall show that, in this case, the wave function conformally transforms under lapse rescalings.